3. what voltage will be generated by the cell that is Constructed by immersing silver wire in AgNo3 Solution of 1.5M and cadmium wire in CdSo4 Solution of 0.75M Cdso4 Solution at 300k, Given Eº Ag = + 0.8 V and E°cd= -0.4 v respectively. Write the cell representation and cell reactions.


1
Expert's answer
2022-03-10T05:22:05-0500

The cell scheme:

А (–) Сd | Cd2+ || Ag| Ag (+) C

cell reaction:

(–) А: Сd – 2e- = Сd2+

(+) C: Ag+ 1e- = Ag


ECd/Cd2+ = –0.40 + (0.059*lg0.75)/2 = –0.404 V

EAg/Ag+ = 0.80 + 0.059*lg1.5 = 0.810 V

emf = 0.810 – (–0.404) = 1.214 V